Summary (Grade 11 NSC Matric Mathematics): Flashcards

📚Flashcards
Summary
Sign up to keep revising.Create a free account to study more flashcards and track your progress.

Practise the cards

10 cards from this deck

Show

Outcome in probability

A single result from an experiment

Sample space

All possible outcomes of an experiment

Event in probability

Collection of one or more outcomes

Range of probability values

Between 0 and 1 (inclusive)

Mutually exclusive events

Events that cannot happen at the same time

Complementary events

Two mutually exclusive events covering all outcomes

Independent events

Events where one doesn't affect the other's probability

Addition rule formula

P(A or B)=P(A)+P(B)−P(A and B)P(A \text{ or } B) = P(A) + P(B) - P(A \text{ and } B)

Complementary rule formula

P(not A)=1−P(A)P(\text{not } A) = 1 - P(A)

Independence rule formula

P(A and B)=P(A)×P(B)P(A \text{ and } B) = P(A) \times P(B)

Join 100,000+ NSC Matric students studying Flashcards with us.

Select your subjects, and get access to A+ resources today.